In a test of the effectiveness of garlic for lowering​ cholesterol, 43 subjects were treated with garlic in a processed tablet form. Cholesterol levels were measured before and after the treatment. The changes in their levels of LDL cholesterol​ (in mg/dL) have a mean of 5.4 and a standard deviation of 17.2. Complete parts​ (a) and​ (b) below.
Answer #1

Ans:

a)

b)Test statistic:

t=(5.4-0)/(17.2/sqrt(43))

t=2.059

df=43-1=42

p-value=tdist(2.059,42,1)=0.0229